OpenStudy (anonymous):

Evaluate the following integrals by converting to polar coordinates: ∫∫ (x+y)dA, R={(x,y)|x^2+y^2≥4, x^2+y^2≤16, R x≥0, y≥0} Please explain step by step

OpenStudy (turingtest):

@Janex are you there to discuss this?

OpenStudy (turingtest):

Graph these equations on the xy-plane. You will find it to be the graph of two concentric circles. You will be integrating between their radii. (notice that the smaller circle has the inequality \(r\ge\), while the larger circle has \(r\le\) ) Their polar formulas should become obvious at that point, as well as the bounds on \(r\) and \(\theta\) Using the same substitutions we did for the previous problem\[x=r\cos\theta\]\[y=r\sin\theta\]\[dA=rdrd\theta\]you can convert the integrand